OpenStudy (anonymous):

A new type of pump can drain a certain pool in 10 hours. An older pump can drain the pool in 15 hours. How long will it take both pumps working together to drain the pool?

OpenStudy (anonymous):

25 hours

OpenStudy (anonymous):

No:(

OpenStudy (anonymous):

are there answer choices?

OpenStudy (anonymous):

Is it Multiple choice?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

also, what is the lesson about

OpenStudy (anonymous):

what concept are they teaching

OpenStudy (anonymous):

Solve the following for x:\[\frac{1}{10}+\frac{1}{15}=\frac{1}{x} \]x = 6